Circular No. 8723

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF PLUTO Further to IAUC 8648, the IAU Working Group for Planetary System

Nomenclature has approved the following new designations and names of satellites of Pluto:

Pluto II Nix  = S/2005 P 2
Pluto III Hydra = S/2005 P 1

SUPERNOVAE 2006db, 2006dc, 2006dd Three additional supernovae have been discovered on unfiltered CCD

images: 2006db by R. Quimby and P. Mondol (cf. IAUC 8622), 2006dc by W. Li (LOSS/KAIT; cf. IAUC 8721), and 2006dd by L. A. G. Monard (cf. IAUC 8709).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006db June 16.18 11^{h}55^{m}38.^{s}69 +44^{o}23′01′′.4 17.2 4′′.6 E, 0′′.5 N
2006dc June 18.36 16 16 04.04 −22 37 16.7 17.8 6′′.3 E, 14′′.6 N
2006dd June 19.17 3 22 41.62 −37 12 13.0 15.0 0′′.3 W, 16′′ N

Additional approximate magnitudes for 2006db: 2005 Mar. 4 and Apr. 8, [18.4 (co-addition of images); 2006 June 17.17 UT, 16.8. Additional KAIT magnitudes for 2006dc in IC 4596: May 20.44, [19.0; 30.42, 18.0; June 19.29, 17.7. Additional approximate magnitudes by Monard for 2006dd in NGC 1316: June 2.153, [17.5 (unfiltered); 20.158, B = 14.66, V = 14.41, R = 14.60, I = 14.30. Nothing is visible at the location of 2006dd on the Digitized Sky Survey (limiting red mag 20.5).

S. Blondin et al. report that 2006db and 2006dc are both type-IIn supernovae (CBETs 551, 552). S. Immler et al. report on CBET 554 that photometry of 2006dd are suggestive of its being a young type-II supernova, with V = 14.0 on June 20.71 UT.

SUPERNOVAE 2006ap, 2006cz, 2006da The following recently discovered supernovae have been classified spec-

troscopically by Blondin et al. as type-Ia events; listed are SN designation, announcement IAUC, date of spectroscopy and approximate of age then with respect to maximum light, and the CBET where the details appear: 2006ap, IAUC 8683, Mar. 4.5 UT, +2 weeks, CBET 418; 2006cz, 8721, June 17, +1 day, 550; 2006da, 8722, June 17, +1 week, 550.

2006 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green